# Mixed Near-field and Far-field Target Localization for Low-altitude Economy

### v1

@@ III-B1 Estimation of angle parameter $\beta$
For UAV $k\in\mathcal{K}$, let $\hat{\beta}_{k}$ denote its estimated $\beta$. However, since $\tilde{\mathbf{a}}_{y}(\beta)=\tilde{\mathbf{a}}_{y}(\beta\pm 1)$ and hence $\widetilde{\mathbf{T}}(\beta)=\widetilde{\mathbf{T}}(\beta\pm 1)$, we have $\widetilde{F}(\hat{\beta}_{k})=\widetilde{F}(\hat{\beta}_{k}\pm 1)$. Therefore, for each estimated true parameter $\hat{\beta}_{k}$, there always exists one ambiguous parameter $\hat{\beta}_{k}^{\rm(am)}$, which is given by
$\hat{\beta}_{k}^{(\rm am)}=\left\{\begin{matrix}\hat{\beta}_{k}+1,\leavevmode\nobreak\ \leavevmode\nobreak\ \hat{\beta}_{k}=\sin\theta_{k}\cos\phi_{k}<0,\\ \hat{\beta}_{k}-1,\leavevmode\nobreak\ \leavevmode\nobreak\ \hat{\beta}_{k}=\sin\theta_{k}\cos\phi_{k}>0.\end{matrix}\right.$ (34)
As such, for $K$ UAVs, a total of $2K$ peaks can be found in the spectrum function $\widetilde{F}(\beta)$ including $K$ true angles and $K$ corresponding ambiguous angles, as illustrated in Fig. 3(a). For convenience, the estimated parameters are collectively denoted as $\tilde{\beta}_{u}\in\{\hat{\beta}_{k}^{(\rm am)},\hat{\beta}_{k}\},u=1,2,\cdots,2K$.
@@ III-B2 Estimation for angle parameter $\alpha$
$\hat{\alpha}_{k}^{{(\rm am)}}=\left\{\begin{matrix}\hat{\alpha}_{k}+1,\leavevmode\nobreak\ \leavevmode\nobreak\ \hat{\alpha}_{k}=\sin\theta_{k}\sin\phi_{k}<0,\\ \hat{\alpha}_{k}-1,\leavevmode\nobreak\ \leavevmode\nobreak\ \hat{\alpha}_{k}=\sin\theta_{k}\sin\phi_{k}>0.\end{matrix}\right.$ (36)
@@
For each parameter pair $(\alpha_{g},\beta_{g})\in\mathcal{A},g\in\mathcal{G}\triangleq\{1,2,\cdots,4K\}$, by solving the equations $\sin\theta_{g}\cos\phi_{g}={\alpha}_{g}$ and $\sin\theta_{g}\sin\phi_{g}={\beta}_{g}$, the candidate angle of the UAV can be obtained as
${\theta}_{g}=\arcsin\sqrt{\hat{\alpha}_{g}^{2}+\hat{\beta}_{g}^{2}},\vspace{-5pt}$ (38)
@@
where ${\theta}_{g}$ and ${\phi}_{g}$ denote candidate elevation and azimuth angles of the UAV. As such, we can obtain $4K$ candidate UAV 2D-angles including the true and ambiguous ones for the far-field and near-field UAVs. Although angular ambiguity cannot be removed in the angle estimation, an effective classification method will be presented in the next to distinguish far-field and near-field UAVs, as well as eliminate angular ambiguity.
Figure 4: Spectrum of far-field estimator in (41) proposed in [17]. System parameters are set as $f=10$ GHz and $N_{x}=N_{y}=61$. The elevation and azimuth angles for far-field UAVs are ($\frac{\pi}{3}$, $-\frac{\pi}{4}$, $1000$ m) and ($\frac{\pi}{8}$, $\frac{2\pi}{3}$, $1500$ m), while the locations for near-field UAVs are set as ($\frac{5\pi}{13}$, ${0.23\pi}$, $30$ m) and ($\frac{\pi}{4}$, $\frac{5\pi}{21}$, $40$ m).
III-C UAV Classification, Near-field UAV Range Estimation and Ambiguity Elimination
In this section, given the candidate 2D-angular parameters of UAVs, we jointly classify the mixed-field UAVs, resolve angular ambiguity, and estimate the ranges of UAVs, by capitalizing on distinct spectrum patterns of far-field and near-field UAVs in the range domain.
@@ III-C UAV Classification, Near-field UAV Range Estimation and Ambiguity Elimination
where ${\mathbf{U}}_{s}\in\mathbb{C}^{N\times K}$ and ${\mathbf{U}}_{n}\in\mathbb{C}^{N\times(N-K)}$ denote the signal and noise subspaces, respectively. Moreover, ${\bm{\Sigma}}_{s}\in\mathbb{C}^{K\times K}$ is a diagonal matrix consisting of $K$ largest eigenvalues, while ${\bm{\Sigma}}_{n}\in\mathbb{C}^{(N-K)\times(N-K)}$ is a diagonal matrix containing $(N-K)$ smallest eigenvalues. Before introducing our proposed classification method, we first discuss the issues in existing mixed-field target classification methods [17, 19].
Issues of existing classification method: In [17, 19], the authors propose to obtain the far-field UAV parameter $\beta$ from the spectrum $F(r,\theta,\phi)$ by letting $r\to\infty$
@@
Since the peaks of far-field UAVs in 3D MUSIC spectrum appears at $r\to\infty$ and $\theta=\theta_{k},\phi=\phi_{k},k\in\mathcal{K}_{1}$ [17], the estimator in (42) can exactly estimate $K_{1}$ far-field UAVs. Similar to Section III-B, the 2D-angular MUSIC estimator in (41) can be decoupled by two 1D MUSIC spectrums. For simplicity, we only provide the estimator for parameter $\beta$, given by
@@
Figure 5: Illustration of proposed classification method. System parameters are set as $f=10$ GHz and $N_{x}=N_{y}=61$. The elevation and azimuth angles for far-field UAVs are ($\frac{\pi}{3}$, $-\frac{\pi}{4}$, $1000$ m) and ($\frac{\pi}{8}$, $\frac{2\pi}{3}$, $1500$ m), while the locations for near-field UAVs are set as ($\frac{5\pi}{13}$, ${0.23\pi}$, $50$ m) and ($\frac{\pi}{4}$, $\frac{5\pi}{21}$, $60$ m).
However, when near-field targets approach the Rayleigh distance, there will exit peaks associated with near-field UAVs in the spectrum $\bar{F}(\beta)$, which results in mistake in the far-field versus near-field target classification. This can be intuitively understood, since in this case, the near-field channel is highly correlated to the far-field channel at the same angle, thus making it possible to be wrongly detected as far-field target. For illustration, we plot the spectrum of $\bar{F}(\beta)$ in Fig. 4, where two near-field UAVs are located at the ranges of $30$ m and $40$ m with the Rayleigh distance given by $Z_{R}=108$ m. It is observed that there exit high peaks associated with near-field UAVs in the spectrum of $\bar{F}(\beta)$, since the the correlation $|\mathbf{b}(\infty,\theta_{3},\phi_{3})^{H}\mathbf{b}(r_{3},\theta_{3},\phi_{3})|=0.96$ and $|\mathbf{b}(\infty,\theta_{4},\phi_{4})^{H}\mathbf{b}(r_{4},\theta_{4},\phi_{4})|=0.98$, which are strongly correlated.
Proposed classification method: To address the above issue in existing works on mixed-field classification, we propose to leverage the characteristics of 3D MUSIC spectrum in the range domain to jointly distinguish far-field and near-field UAVs, estimate ranges of UAVs, as well as eliminate the ambiguity in angle estimation obtained in Section (III-B).
To this end, for each candidate UAV 2D-angles, we consider the original covariance matrix $\mathbf{R}_{\bm{\eta}}$ and obtain its MUSIC spectrum in the range domain given the 2D-angles, which is given by
@@
where $g\in\mathcal{G}$ denotes the candidate UAV angle index. Then, in Fig. 5, we numerically show the 3D MUSIC spectrum in the range domain given the 2D-angles of the near-field UAV, far-field UAV, as well as ambiguous 2D-angles. Several important observations are made as follows.
@@
    (Near-field UAV angle) When $\{\theta_{g},\theta_{g}\}$ are the 2D-angles of a near-field UAV, its 3D MUSIC spectrum in the range domain will first increases and then decreases, when the range is smaller than the Rayleigh distance (shown in Fig. 5), which is referred to as Pattern 1. In addition, it achieves the peak value in the spectrum $F_{g}(r,{\theta}_{g},{\phi}_{g})$ at the UAV range, which is smaller than the Rayleigh distance $Z_{\rm R}$. (see Fig. 5).
@@
    (Far-field UAV angle) When $\{\theta_{g},\theta_{g}\}$ are the 2D-angles of a near-field UAV, its 3D MUSIC spectrum in the range domain will monotonically increase when $r\leq Z_{\rm R}$, which is called as Pattern 2. This is expected since the peak value of the spectrum $F_{g}(r,{\theta}_{g},{\phi}_{g})$ is achieved at $r\to\infty$.
@@
The above observations indicate that the three types of estimated UAV 2D-angles exhibit significantly different patterns in the range-domain MUSIC spectrum. This thus leads to the proposed joint UAV classification-localization method, as presented below.
$F_{g}(r,\theta_{g},\phi_{g})\in$ Pattern 1 $(\theta_{g},\phi_{g})$ are 2D-angles of a near-field UAV. $\hat{r}=\arg\max\limits_{r}F_{g}(r,\theta_{g},\phi_{g})$.
$F_{g}(r,\theta_{g},\phi_{g})\in$ Pattern 2 $(\theta_{g},\phi_{g})$ are 2D-angles of a far-field UAV.
$F_{g}(r,\theta_{g},\phi_{g})\in$ Pattern 3 $(\theta_{g},\phi_{g})$ are ambiguous angles.
@@
The computational complexity of proposed mixed-field localization method is dominated by three parts: 1) pre-processing; 2) candidate angle parameter estimation; 3) joint UAV classification, angle ambiguity elimination and near-field UAV range estimation. In particular, the complexity of pre-processing is dominant in (13), which amounts to $\mathcal{O}(LN^{3})$, due to the matrix inverse of $\mathbf{W}$ for $L$ times. For the angle parameter estimation, the dominant complexity mainly involves in the eigenvalue decomposition in (26) that is $\mathcal{O}\Big{(}\frac{(N_{x}+1)^{3}(N_{y}+1)^{3}}{64}\Big{)}$ and the grid searches for parameter $\beta_{k}$ and $\alpha_{k}$ which are $\mathcal{O}\Big{(}\ell_{y}\frac{(N_{x}+1)^{2}}{16}+\ell_{y}\frac{(N_{x}+1)^{3}}{8}\Big{)}$ and $\mathcal{O}\Big{(}2K\ell_{x}N^{2}\Big{)}$, where $\ell_{x}$ and $\ell_{y}$ denote the number of search grids for parameters $\alpha$ and $\beta$, respectively [18]. Then, for the joint UAV classification, near-field UAV range estimation, and ambiguity elimination, the dominant complexity is contributed by the grid-search in (43), given by $\mathcal{O}(4K\ell_{d}N^{2})$, where $\ell_{d}$ represents the number of search grids for distance. Hence, the overall computational complexity of the proposed localization method is $\mathcal{O}(K\ell_{x}N^{2}+\ell_{y}N_{x}^{3}+K\ell_{d}N^{2}+LN^{3})$. As compared to the proposed methods, the 3D MUSIC algorithm incurs a heavy computational complexity of $\mathcal{O}(\ell_{x}\ell_{y}\ell_{d}N^{2}+LN^{3})$, since the number of searched grids satisfy $\ell_{x},\ell_{y},\ell_{d}\gg N$. Therefore, the proposed decoupled 3D MUSIC-based localization method achieves significant reduction in computational complexity as compared to the 3D MUSIC algorithm.
@@ III-D Extensions and Discussions
Low-complexity MUSIC algorithm: When the received signal power of far-field UAVs is comparable to that near-field UAVs, we can employ the 2D DFT-based method [15] to coarsely estimate the angle parameters for both far-field and near field UAVs. As such, we only need to perform grid search in (33) and (35) around coarse angles estimated by the DFT-method, which significantly reduces the computational complexity in grid search.
@@
$[\mathbf{M}]_{i,j}\!\!=\!\!\sum_{k=1}^{K}g_{k}e^{\jmath 2d_{0}(i\alpha_{k}\!+\!j\beta_{k})}e^{-\jmath 2d_{0}(\frac{N_{x}+1}{2}\alpha_{k}\!+\!\frac{N_{y}+1}{2}\beta_{k})},$ (45)
where $i=1,2,\cdots,N_{x}$ and $j=1,2,\cdots,N_{y}$. Then, we define two DFT matrices $\mathbf{D}_{x}\in\mathbb{C}^{N_{x}\times N_{x}}$ and $\mathbf{D}_{y}\in\mathbb{C}^{N_{y}\times N_{y}}$, which are given by $[\mathbf{D}_{x}]_{i_{x},j_{x}}=\frac{1}{N_{x}}e^{-\jmath 2\pi\frac{(i_{x}-1)(j_{x}-1)}{N_{x}}},i_{x},j_{x}=1,2,\cdots,N_{x}$ and $[\mathbf{D}_{y}]_{i_{y},j_{y}}=\frac{1}{N_{y}}e^{-\jmath 2\pi\frac{(i_{y}-1)(j_{y}-1)}{N_{y}}},i_{y},j_{y}=1,2,\cdots,N_{y}.$ As such, the normalized 2D-DFT of the constructed matrix $\mathbf{M}$ can be calculated as
[…]
@@
It can be observed that when $(\alpha_{k}-{i}/{N_{x}})\pi=\ell_{x}\pi,\forall\ell_{x}\in\mathbb{Z}$ and $(\beta_{k}-{j}/{N_{y}})\pi=\ell_{y}\pi,\forall\ell_{y}\in\mathbb{Z}$ (i.e., $i=N_{x}(\alpha_{k}-1)$ and $j=N_{y}(\beta_{k}-1)$), $\big{|}\mathbf{M}_{\rm DFT}[{i,j}]\big{|}$ forms a peak with a height $|g_{k}|$. Hence, by denoting ($\hat{i}_{k}$, $\hat{j}_{k}$) as an integer pair where $\big{|}[\mathbf{M}_{\rm DFT}]_{{\hat{i}_{k},\hat{j}_{k}}}\big{|}$ reaches a peak value, the parameter estimation for the $k$-th UAV is given by $\hat{\alpha}_{k}^{(1)}=\frac{\hat{i}_{k}}{N_{x}},\hat{\alpha}_{k}^{(2)}=\frac{\hat{i}_{k}}{N_{x}}-1$, and $\hat{\beta}_{k}^{(1)}=\frac{\hat{j}_{k}}{N_{y}},\hat{\beta}_{k}^{(2)}=\frac{\hat{j}_{k}}{N_{y}}-1$, where $\hat{\alpha}_{k}^{(1)}$ and $\hat{\alpha}_{k}^{(2)}$ are estimated parameter angles for the $k$-th UAV, corresponding to the true and ambiguous angle parameters, respectively. Similarly, $\hat{\beta}_{k}^{(1)}$ and $\hat{\beta}_{k}^{(2)}$ follow the same pattern as $\hat{\alpha}_{k}^{(1)}$ and $\hat{\alpha}_{k}^{(2)}$. As such, we obtain the coarse angle parameters $\{\hat{\alpha}_{k}^{(1)},\hat{\alpha}_{k}^{(2)},\hat{\beta}_{k}^{(1)},\hat{\beta}_{k}^{(2)},\forall k\in\mathcal{K}\}$ with a resolution $1/N_{x}$ and $1/N_{y}$ for the parameter $\alpha$ and $\beta$ estimation, respectively. To facilitate understanding, in Fig. 6, we numerically plot an example of 2D-DFT of the constructed matrix $\mathbf{M}$ with $N_{x}=N_{y}=101$.

### v2

@@ III-B1 Estimation of angle parameter $\beta$
For target $k\in\mathcal{K}$, let $\hat{\beta}_{k}$ denote its estimated $\beta$. However, since $\tilde{\mathbf{a}}_{y}(\beta)=\tilde{\mathbf{a}}_{y}(\beta\pm 1)$ and hence $\widetilde{\mathbf{T}}(\beta)=\widetilde{\mathbf{T}}(\beta\pm 1)$, we have $\widetilde{F}(\hat{\beta}_{k})=\widetilde{F}(\hat{\beta}_{k}\pm 1)$. Therefore, for each estimated true parameter $\hat{\beta}_{k}$, there always exists one ambiguous parameter $\hat{\beta}_{k}^{\rm(am)}$, which is given by
$\hat{\beta}_{k}^{(\rm am)}=\left\{\begin{matrix}\hat{\beta}_{k}+1,\penalty 10000\ \penalty 10000\ \hat{\beta}_{k}=\sin\theta_{k}\cos\phi_{k}<0,\\ \hat{\beta}_{k}-1,\penalty 10000\ \penalty 10000\ \hat{\beta}_{k}=\sin\theta_{k}\cos\phi_{k}>0.\end{matrix}\right$. (34)
As such, for $K$ targets, a total of $2K$ peaks can be found in the spectrum function $\widetilde{F}(\beta)$ including $K$ true angles and $K$ corresponding ambiguous angles, as illustrated in Fig. 3(a). For convenience, the estimated parameters are collectively denoted as $\tilde{\beta}_{u}\in\{\hat{\beta}_{k}^{(\rm am)},\hat{\beta}_{k}\},u=1,2,\cdots,2K$.
@@ III-B2 Estimation for angle parameter $\alpha$
$\hat{\alpha}_{k}^{{(\rm am)}}=\left\{\begin{matrix}\hat{\alpha}_{k}+1,\penalty 10000\ \penalty 10000\ \hat{\alpha}_{k}=\sin\theta_{k}\sin\phi_{k}<0,\\ \hat{\alpha}_{k}-1,\penalty 10000\ \penalty 10000\ \hat{\alpha}_{k}=\sin\theta_{k}\sin\phi_{k}>0.\end{matrix}\right$. (36)
@@
Figure 4: Spectrum of far-field estimator in (41) proposed in [6]. System parameters are set as $f=10$ GHz and $N_{x}=N_{y}=61$. The elevation and azimuth angles for far-field targets are ($\frac{\pi}{3}$, $\frac{\pi}{4}$, $1000$ m) and ($\frac{\pi}{8}$, $\frac{2\pi}{3}$, $1500$ m), while the locations for near-field targets are set as ($\frac{5\pi}{13}$, $0.23\pi$}, $30$ m) and ($\frac{\pi}{4}$, $\frac{5\pi}{21}$, $40$ m).
@@
For each parameter pair $(\alpha_{g},\beta_{g})\in\mathcal{A},g\in\mathcal{G}\triangleq\{1,2,\cdots,4K\}$, by solving the equations $\sin\theta_{g}\cos\phi_{g}={\alpha}_{g}$ and $\sin\theta_{g}\sin\phi_{g}={\beta}_{g}$, the candidate angle of the target can be obtained as
${\theta}_{g}=\arcsin\sqrt{\hat{\alpha}_{g}^{2}+\hat{\beta}_{g}^{2}},\vskip-5.0pt$ (38)
@@
where ${\theta}_{g}$ and ${\phi}_{g}$ denote candidate elevation and azimuth angles of the target. As such, we can obtain $4K$ candidate 2D angles including the true and ambiguous ones for the far-field and near-field targets. Although angular ambiguity cannot be removed in the angle estimation, an effective classification method will be presented in the next to distinguish far-field and near-field targets, as well as eliminate angular ambiguity.
III-C Target Classification, Near-field Target Range Estimation and Ambiguity Elimination
In this section, given the candidate 2D-angular parameters of targets, we jointly classify the mixed-field targets, resolve angular ambiguity, and estimate the ranges of targets, by capitalizing on distinct spectrum patterns of far-field and near-field targets in the range domain.
@@ III-C Target Classification, Near-field Target Range Estimation and Ambiguity Elimination
where ${\mathbf{U}}_{s}\in\mathbb{C}^{N\times K}$ and ${\mathbf{U}}_{n}\in\mathbb{C}^{N\times(N-K)}$ denote the signal and noise subspaces, respectively. Moreover, ${\bm{\Sigma}}_{s}\in\mathbb{C}^{K\times K}$ is a diagonal matrix consisting of $K$ largest eigenvalues, while ${\bm{\Sigma}}_{n}\in\mathbb{C}^{(N-K)\times(N-K)}$ is a diagonal matrix containing (NK)(N-K) smallest eigenvalues. Before introducing our proposed classification method, we first discuss the issues in existing mixed-field target classification methods [6, 10].
Issues of existing classification method: In [6, 10], the authors propose to obtain the far-field target parameter $\beta$ from the spectrum $F(r,\theta,\phi)$ by letting $r\to\infty$
@@
Since the peaks of far-field targets in 3D MUSIC spectrum appears at $r\to\infty$ and $\theta=\theta_{k},\phi=\phi_{k},k\in\mathcal{K}_{1}$ [6], the estimator in (42) can exactly estimate $K_{1}$ far-field targets. Similar to Section III-B, the 2D-angular MUSIC estimator in (41) can be decoupled by two 1D MUSIC spectrums. For simplicity, we only provide the estimator for parameter $\beta$, given by
@@
Figure 5: Illustration of proposed classification method. System parameters are set as $f=10$ GHz and $N_{x}=N_{y}=61$. The ranges of the two near-field targets are $50$ and $60$ m.
Note that, when near-field targets approach the Rayleigh distance, there will exit peaks associated with near-field targets in the spectrum $\bar{F}(\beta)$, which results in mistake in the far-field versus near-field target classification. This can be intuitively understood, since in this case, the near-field channel is highly correlated to the far-field channel at the same angle, thus making it possible to be wrongly detected as far-field target. For illustration, we plot the spectrum of $\bar{F}(\beta)$ in Fig. 4, where two near-field targets are located at the ranges of $30$ m and $40$ m with the Rayleigh distance given by $Z_{R}=108$ m. It is observed that there exit high peaks associated with near-field targets in the spectrum of $\bar{F}(\beta)$, since the the correlation $|\mathbf{b}(\infty,\theta_{3},\phi_{3})^{H}\mathbf{b}(r_{3},\theta_{3},\phi_{3})|=0.96$ and $|\mathbf{b}(\infty,\theta_{4},\phi_{4})^{H}\mathbf{b}(r_{4},\theta_{4},\phi_{4})|=0.98$, which are strongly correlated.
Proposed  classification method  under  half-wavelength inter-antenna spacing: To address the above issue in existing works on mixed-field classification, we propose to leverage the characteristics of 3D MUSIC spectrum in the range domain to jointly distinguish far-field and near-field targets, estimate ranges of targets, as well as eliminate the ambiguity in angle estimation obtained in Section (III-B).
To this end, for each candidate target 2D angles, we consider the original covariance matrix $\mathbf{R}_{\bm{\eta}}$ and obtain its MUSIC spectrum in the range domain given the 2D angles, which is given by
@@
where $g\in\mathcal{G}$ denotes the candidate target angle index. Then, in Fig. 5, we numerically show the 3D MUSIC spectrum in the range domain given the 2D angles of the near-field target, far-field target, as well as ambiguous 2D angles. Several important observations are made as follows.
@@
    (Near-field target angle) When $\{\theta_{g},\theta_{g}\}$ are the 2D angles of a near-field target, its 3D MUSIC spectrum in the range domain will first increases and then decreases, when the range is smaller than the Rayleigh distance (shown in Fig. 5), which is referred to as Pattern 1. In addition, it achieves the peak value in the spectrum $F_{g}(r,{\theta}_{g},{\phi}_{g})$ at the target range, which is smaller than the Rayleigh distance $Z_{\rm R}$. (see Fig. 5).
@@
    (Far-field target angle) When $\{\theta_{g},\theta_{g}\}$ are the 2D angles of a near-field target, its 3D MUSIC spectrum in the range domain will monotonically increase when $r\leq Z_{\rm R}$, which is called as Pattern 2. This is expected since the peak value of the spectrum $F_{g}(r,{\theta}_{g},{\phi}_{g})$ is achieved at $r\to\infty$.
@@
The above observations indicate that the three types of estimated 2D angles exhibit significantly different patterns in the range-domain MUSIC spectrum. This thus leads to the proposed joint target classification-localization method, as presented below.
$F_{g}(r,\theta_{g},\phi_{g})\in$ Pattern 1 $(\theta_{g},\phi_{g})$ are 2D angles of a near-field target. $\hat{r}=\arg\max\limits_{r}F_{g}(r,\theta_{g},\phi_{g})$. $F_{g}(r,\theta_{g},\phi_{g})\in$ Pattern 2 $(\theta_{g},\phi_{g})$ are 2D angles of a far-field target. $F_{g}(r,\theta_{g},\phi_{g})\in$ Pattern 3 $(\theta_{g},\phi_{g})$ are ambiguous angles.
@@
The computational complexity of proposed mixed-field localization method is dominated by three parts: 1) pre-processing; 2) candidate angle parameter estimation; 3) joint target classification, angle ambiguity elimination and near-field target range estimation. In particular, the complexity of pre-processing is dominant in (13), which amounts to $\mathcal{O}(LN^{3})$, due to the matrix inverse of $\mathbf{W}$ for $L$ times. For the angle parameter estimation, the dominant complexity mainly involves in the eigenvalue decomposition in (26) that is $\mathcal{O}\Big(\frac{(N_{x}+1)^{3}(N_{y}+1)^{3}}{64}\Big)$ and the grid searches for parameter $\beta_{k}$ and $\alpha_{k}$ which are $\mathcal{O}\Big(\ell_{y}\frac{(N_{x}+1)^{2}}{16}+\ell_{y}\frac{(N_{x}+1)^{3}}{8}\Big)$ and $\mathcal{O}\Big(2K\ell_{x}N^{2}\Big)$, where $\ell_{x}$ and $\ell_{y}$ denote the number of search grids for parameters $\alpha$ and $\beta$, respectively [8]. Then, for the joint target classification, near-field target range estimation, and ambiguity elimination, the dominant complexity is contributed by the grid-search in (43), given by $\mathcal{O}(4K\ell_{d}N^{2})$, where $\ell_{d}$ represents the number of search grids for distance. Hence, the overall computational complexity of the proposed localization method is $\mathcal{O}(K\ell_{x}N^{2}+\ell_{y}N_{x}^{3}+K\ell_{d}N^{2}+LN^{3})$. As compared to the proposed methods, the 3D MUSIC algorithm incurs a heavy computational complexity of $\mathcal{O}(\ell_{x}\ell_{y}\ell_{d}N^{2}+LN^{3})$, since the number of searched grids satisfy $\ell_{x},\ell_{y},\ell_{d}\gg N$. Therefore, the proposed decoupled 3D MUSIC-based localization method achieves significant reduction in computational complexity as compared to the 3D MUSIC algorithm.
@@ III-D Extensions and Discussions
Low-complexity MUSIC algorithm: When the received signal power of far-field targets is comparable to that near-field targets, we can employ the 2D DFT-based method [4] to coarsely estimate the angle parameters for both far-field and near field targets. As such, we only need to perform grid search in (33) and (35) around coarse angles estimated by the DFT-method, which significantly reduces the computational complexity in grid search.
@@
$\mathbf{M}]_{i,j}\!\!=\!\!\sum_{k=1}^{K}g_{k}e^{\jmath 2d_{0}(i\alpha_{k}\!+\!j\beta_{k})}e^{-\jmath 2d_{0}(\frac{N_{x}+1}{2}\alpha_{k}\!+\!\frac{N_{y}+1}{2}\beta_{k})},$ (45)
where $i=1,2,\cdots,N_{x}$ and $j=1,2,\cdots,N_{y}$. Then, we define two DFT matrices $\mathbf{D}_{x}\in\mathbb{C}^{N_{x}\times N_{x}}$ and $\mathbf{D}_{y}\in\mathbb{C}^{N_{y}\times N_{y}}$, which are given by $\mathbf{D}_{x}]_{i_{x},j_{x}}=\frac{1}{N_{x}}e^{-\jmath 2\pi\frac{(i_{x}-1)(j_{x}-1)}{N_{x}}},i_{x},j_{x}=1,2,\cdots,N_{x}$ and $\mathbf{D}_{y}]_{i_{y},j_{y}}=\frac{1}{N_{y}}e^{-\jmath 2\pi\frac{(i_{y}-1)(j_{y}-1)}{N_{y}}},i_{y},j_{y}=1,2,\cdots,N_{y}.$ As such, the normalized 2D-DFT of the constructed matrix $\mathbf{M}$ can be calculated as
$\big|$ $[\mathbf{M}_{\rm DFT}]_{{i,j}}\big|=\big|[\mathbf{D}_{x}\mathbf{M}\mathbf{D}_{y}]_{{i,j}}\big|$
@@
It can be observed that when $(\alpha_{k}-{i}/{N_{x}})\pi=\ell_{x}\pi,\forall\ell_{x}\in\mathbb{Z}$ and $(\beta_{k}-{j}/{N_{y}})\pi=\ell_{y}\pi,\forall\ell_{y}\in\mathbb{Z}$ (i.e., $i=N_{x}(\alpha_{k}-1)$ and $j=N_{y}(\beta_{k}-1)$), $\big|\mathbf{M}_{\rm DFT}[{i,j}]\big|$ forms a peak with a height $|g_{k}|$. Hence, by denoting ($\hat{i}_{k}$, $\hat{j}_{k}$) as an integer pair where $\big|[\mathbf{M}_{\rm DFT}]_{{\hat{i}_{k},\hat{j}_{k}}}\big|$ reaches a peak value, the parameter estimation for the $k$-th target is given by $\hat{\alpha}_{k}^{(1)}=\frac{\hat{i}_{k}}{N_{x}},\hat{\alpha}_{k}^{(2)}=\frac{\hat{i}_{k}}{N_{x}}-1$, and $\hat{\beta}_{k}^{(1)}=\frac{\hat{j}_{k}}{N_{y}},\hat{\beta}_{k}^{(2)}=\frac{\hat{j}_{k}}{N_{y}}-1$, where $\hat{\alpha}_{k}^{(1)}$ and $\hat{\alpha}_{k}^{(2)}$ are estimated parameter angles for the $k$-th target, corresponding to the true and ambiguous angle parameters, respectively. Similarly, $\hat{\beta}_{k}^{(1)}$ and $\hat{\beta}_{k}^{(2)}$ follow the same pattern as $\hat{\alpha}_{k}^{(1)}$ and $\hat{\alpha}_{k}^{(2)}$. As such, we obtain the coarse angle parameters $\{\hat{\alpha}_{k}^{(1)},\hat{\alpha}_{k}^{(2)},\hat{\beta}_{k}^{(1)},\hat{\beta}_{k}^{(2)},\forall k\in\mathcal{K}\}$ with a resolution $1/N_{x}$ and $1/N_{y}$ for the parameter $\alpha$ and $\beta$ estimation, respectively. To facilitate understanding, in Fig. 6, we numerically plot an example of 2D-DFT of the constructed matrix $\mathbf{M}$ with $N_{x}=N_{y}=101$.